# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ VIII Conclusion
      Any element of $\mathcal{T}_{g_{0}}$ can be expressed as an element of a given $\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$, for some $\tilde{i}\in\mathbb{N}$, or as a converging sequence of such elements, i.e. $\mathcal{T}_{g_{0}}\subset\overline{\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}}$ where $\overline{\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}}$ is the closure in $\mathcal{H}$ of $\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}$.
    Let us start by showing i). Each element of $\mathcal{T}_{\bm{\rho}_{0},i},\forall i\in\mathbb{N}$ is of the form $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i},\mathbf{c}_{i}\in\mathbb{R}^{r_{i}}$ where:
    $\mathbf{s}_{\bm{\rho}_{0},i}=\nabla_{\bm{\rho}}\ln\left[s_{0}^{-m/2}|\mathbf{V}_{S,0}|^{-1/2}\tau_{\bm{\rho}_{0,i}}\left(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})\right)\right].$ (167)
    Consequently, $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}$ is clearly an element of $\mathcal{T}_{g_{0}}$ thanks to the linearity of the expectation operator and to the fact that $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\}=0$ and $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]^{2}_{j}\}<+\infty$, for $j=1,\ldots,r_{i}$ from usual properties of the score vectors (under the regularity conditions discussed in [17, Sects. 6.2, 6.3]). Then $\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\in\mathcal{H}$ for $j=1,\ldots,r_{i}$ and it is measurable w.r.t. $Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})$, then $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}\in\mathcal{T}_{g_{0}},\forall i\in\mathbb{N}$.
    Let us now move to ii). Let star by choosing $r_{\tilde{i}}$ elements $\{\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\}_{j=1}^{r_{\tilde{i}}}\in\mathcal{T}_{g_{0}}$ such that $\tilde{h}_{j}$ are bounded functions. Then, as parametric sub-model of the form in (162), we may choose the following one:
    $\mathcal{P}_{\tau_{\bm{\rho},\tilde{i}}}=\left\{p_{X}(\mathbf{x}|\bm{\rho})=p_{0}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\left[1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}\right]\right\}.$ (168)
@@
    $1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}\geq 0,\forall\mathbf{x}\in\mathcal{X},$ (169)
@@
    $\begin{split}\int p_{X}(\mathbf{x}|\bm{\rho})d\mathbf{x}&=\int p_{0}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))d\mathbf{x}+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\int\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))p_{0}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))d\mathbf{x}\\ &=1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}{E_{0}\{\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\}}=1+0=1,\end{split}$ (170)
@@
    since $\tilde{h}_{j}\in\mathcal{H}$, then $p_{X}(\mathbf{x}|\bm{\rho})$ is a proper pdf and consequently $\mathcal{P}_{\tau_{\bm{\rho},\tilde{i}}}$ is a proper parametric sub-model that satisfies the condition C0, C1 and C2. Now, a score vector of this specific parametric sub-model is of the form:
@@
    $\begin{split}\mathbf{s}_{\bm{\rho}_{0},\tilde{i}}&=\nabla_{\bm{\rho}}\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho}_{0},i})=\\ &=\left.\nabla_{\bm{\rho}}\ln\left[p_{0}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\left(1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}\right)\right]\right|_{\bm{\rho}=\bm{\rho}_{0}}\\ &=\left.\nabla_{\bm{\rho}}\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}\right|_{\bm{\rho}=\bm{\rho}_{0}}=(\tilde{h}_{1},\ldots,\tilde{h}_{r_{\tilde{i}}})^{T}.\end{split}$ (171)
@@
    Consequently, since, as said before, any element of $\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$ is of the form $\mathbf{c}^{T}\mathbf{s}_{\bm{\rho}_{0},\tilde{i}}$ for some $\mathbf{c}\in\mathbb{R}^{r_{\tilde{i}}}$, we just need to choose $\mathbf{c}=\mathbf{e}_{j,r_{\tilde{i}}}$ to prove that $\mathcal{T}_{g_{0}}\ni\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\in\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$, where $\tilde{h}_{j}$ is a bounded function. The proof is completed by noticing that the set of bounded functions is dense in $\mathcal{H}$ and consequently, any element $h\in\mathcal{T}_{g_{0}}$ can be obtained as a converging sequence of bounded functions. This allow us to state that $\mathcal{T}_{g_{0}}\subset\overline{\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}}$ and conclude the proof. ∎
  8. .8
@@
    Explicit expressions of the CRBs in the models $\mathcal{P}_{\bm{\eta},g_{0}}$ and $\mathcal{P}_{\bm{\nu},g_{0}}$. Let us first evaluate the CRB on $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}\in\Omega$ in (41) as the inverse of the related FIM $\mathbf{I}_{\bm{\nu}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\nu}_{0}}\mathbf{s}_{\bm{\nu}_{0}}^{T}\}$. To this end, we note that the two components of the related score vector $\mathbf{s}_{\bm{\nu}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T})^{T}$ have been already introduced in (56) and (57). Consequently, from standard calculation and by using the independence between $\mathcal{Q}$ and $\mathbf{u}$ (along with the properties of $\mathbf{u}$), the FIM for $\bm{\nu}_{0}\in\Omega$ in $\mathcal{P}_{\bm{\nu},g_{0}}$ is given by:
@@
    $\mathbf{I}_{\bm{\mu}_{0}}\triangleq E_{0}\left\{\mathbf{s}_{\bm{\mu}_{0}}\mathbf{s}_{\bm{\mu}_{0}}^{T}\right\}=\beta(g_{0})\bm{\Sigma}_{0}^{-1},$ (173)
@@
    $\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\triangleq E_{0}\left\{\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T}\right\}=\mathbf{D}_{m}^{T}\left[\frac{1}{2}\alpha(g_{0})(\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1})+\frac{1}{4}({\alpha(g_{0})-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}\right]\mathbf{D}_{m}.$ (175)
    Then the CRB on $\bm{\nu}_{0}$ in (41) can be derived from the block diagonal structured FIM in (172) as:
@@
    $\mathrm{CRB}(\bm{\mu}_{0})\triangleq\mathbf{I}_{\bm{\mu}_{0}}^{-1}=\beta(g_{0})^{-1}\bm{\Sigma}_{0}\ \ \mbox{and}\ \ \mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))\triangleq\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{-1}.$ (177)
@@
    $\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]^{-1}=2\alpha^{-1}(g_{0})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{2\alpha^{-1}(g_{0})(\alpha(g_{0})-1)}{(m+2)\alpha(g_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}.$ (178)
    Then using $\mathbf{D}_{m}\mathbf{D}_{m}^{\#}=\frac{1}{2}(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})$ and $\mathbf{K}_{m}(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})=(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})\mathbf{K}_{m}$ [23, Ch. 3], we straightforwardly get
    $\mathbf{D}_{m}\mathbf{D}_{m}^{\#}[\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]^{-1}\mathbf{D}_{m}^{\#T}\mathbf{D}^{T}_{m}=\alpha^{-1}(g_{0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{2\alpha^{-1}(g_{0})(\alpha(g_{0})-1)}{(m+2)\alpha(g_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}.$ (179)
    Finally, using $\mathbf{K}_{m}\mathbf{D}_{m}=\mathbf{D}_{m}$, $\mathbf{D}_{m}^{\#}\mathbf{D}_{m}=\mathbf{I}_{m(m+1)/2}$, [23, Ch. 3], it is easy to check that
@@
    and thus, since $\mathbf{D}_{m}^{\#}{\mathbf{K}}_{m}=\mathbf{D}_{m}^{\#}$ [23, Ch. 3]
    $\mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))=2\alpha^{-1}(g_{0})\mathbf{D}_{m}^{\#}\left[(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{(\alpha(g_{0})-1)}{(m+2)\alpha(g_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}\right]\mathbf{D}_{m}^{\#T}.$ (181)
@@
    $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})=\alpha^{-1}(g_{0})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},$ (190)

### v3

@@ VIII Conclusion
      Any element of $\mathcal{T}^{u}_{g_{0}}$ can be expressed as an element of a given $\mathcal{T}^{u}_{\tilde{i},\bm{\rho}_{0}}$, for some $\tilde{i}\in\mathbb{N}$, or as a converging sequence of such elements.
    Let us start by showing i). Each element of $\mathcal{T}^{u}_{i,\bm{\rho}_{0}},\forall i\in\mathbb{N}$ is of the form $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}$, where $\mathbf{c}_{i}\in\mathbb{R}^{r_{i}}$ and:
    $\color[rgb]{0,0,0}\mathbf{s}_{\bm{\rho}_{0},i}=\nabla_{\bm{\rho}}\ln\left[|\mathbf{V}_{S,0}|^{-1/2}\tau_{i,\bm{\rho}_{0}}\left(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x})\right)\right]$}. (185)
    Consequently, $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}$ is an element of $\mathcal{T}^{u}_{g_{0}}$, since, by inverting the order of differentiation and integration (under the regularity conditions discussed in [19, Sects. 6.2, 6.3]), it is immediate to verify that $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\}=0$, for $j=1,\ldots,r_{i}$. Then $\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\in\mathcal{H}$ for $j=1,\ldots,r_{i}$ and it is measurable w.r.t. $Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x})$, then $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}\in\mathcal{T}^{u}_{g_{0}},\forall i\in\mathbb{N}$.
    Let us now move to ii). Let start by choosing $r_{\tilde{i}}$ elements $\{\tilde{h}_{j}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x}))\}_{j=1}^{r_{\tilde{i}}}\in\mathcal{T}^{u}_{g_{0}}$ such that $\tilde{h}_{j}$ are bounded functions. Then, as parametric sub-model of the form in (180), we may choose the following one:
    $\color[rgb]{0,0,0}\mathcal{P}^{u}_{\tau_{\bm{\rho},\tilde{i}}}=\left\{p_{X}(\mathbf{x}|\bm{\rho})=p_{0}(\mathbf{x})\left[1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x}))\right]\right\},$} (186)
@@
    $1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}{\color[rgb]{0,0,0}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x}))}\geq 0,\forall\mathbf{x}\in\mathcal{X},$ (187)
@@
    $\begin{split}\int p_{X}(\mathbf{x}|\bm{\rho})d\mathbf{x}&={\color[rgb]{0,0,0}\int p_{0}(\mathbf{x})d\mathbf{x}+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\int\tilde{h}_{j}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x}))p_{0}(\mathbf{x})d\mathbf{x}}\\ &{\color[rgb]{0,0,0}=1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}{E_{0}\{\tilde{h}_{j}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x}))\}}}=1+0=1,\end{split}$ (188)
@@
    since $\tilde{h}_{j}\in\mathcal{H}$, then $p_{X}(\mathbf{x}|\bm{\rho})$ is a proper pdf and consequently $\mathcal{P}_{\tau_{\bm{\rho},\tilde{i}}}$ is a proper parametric sub-model that satisfies the conditions C1 and C2. Specifically, C2 is verified for $\bm{\rho}_{0}=\mathbf{0}$. Now, a score vector of this specific parametric sub-model is of the form:
@@
    $\begin{split}\mathbf{s}_{\bm{\rho}_{0},\tilde{i}}&={\color[rgb]{0,0,0}\nabla_{\bm{\rho}}\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},\tau_{i,\bm{\rho}_{0}})}=\\ &={\color[rgb]{0,0,0}\left.\nabla_{\bm{\rho}}\ln\left[p_{0}(\mathbf{x})\left(1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x})\right)\right]\right|_{\bm{\rho}=\mathbf{0}}}\\ &={\color[rgb]{0,0,0}\left.\nabla_{\bm{\rho}}\ln\left[1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x})\right]\right|_{\bm{\rho}=\mathbf{0}}=(\tilde{h}_{1},\ldots,\tilde{h}_{r_{\tilde{i}}})^{T}}.\end{split}$ (189)
@@
    Consequently, since, as said before, any element of $\mathcal{T}^{u}_{\tilde{i},\bm{\rho}_{0}}$ is of the form $\mathbf{c}^{T}\mathbf{s}_{\bm{\rho}_{0},\tilde{i}}$ for some $\mathbf{c}\in\mathbb{R}^{r_{\tilde{i}}}$, we just need to choose $\mathbf{c}=\mathbf{e}_{j,r_{\tilde{i}}}$ to prove that $\mathcal{T}^{u}_{g_{0}}\ni{\color[rgb]{0,0,0}\tilde{h}_{j}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x}))}\in\mathcal{T}^{u}_{\tilde{i},\bm{\rho}_{0}}$, where $\tilde{h}_{j}$ is a bounded function. The proof is completed by noticing that the set of bounded functions is dense in $\mathcal{H}$ and consequently, any element $h\in\mathcal{T}^{u}_{g_{0}}$ can be obtained as a converging sequence of bounded functions. This concludes the proof of the result in (77). We conclude this part by noticing that, $\Pi(h|\mathcal{T}^{u}_{g_{0}})=E\{h|\mathcal{Q}\},\;\forall h\in\mathcal{H}$ [27, Ch. 23, Def. 4].
    Let us now move to the constrained non-parametric sub-model $\mathcal{P}^{c}_{\bar{g}}$ in (46). To derive the relevant tangent space $\mathcal{T}^{c}_{\bar{g}_{0}}$, we can reuse the previous proof except for the fact that we need to impose such constraint to the class of parametric sub-model given in (186). Specifically, since form (32) $\mathcal{Q}=_{d}Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x})$, we want that:
    $\begin{split}m=&E\{Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x})\}=\int Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x})p_{0}(\mathbf{x})\left[1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}(Q_{\bm{\mu}_{0},\mathbf{V}_{S,0}}(\mathbf{x}))\right]d\mathbf{x}\\ &=E\{\mathcal{Q}\}+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}E\{\mathcal{Q}\tilde{h}_{j}(\mathcal{Q})\},\end{split}$ (190)
    for all the possible $\{\rho_{j}\}_{j=1}^{r_{\tilde{i}}}$ and then we should impose that $E\{\mathcal{Q}\tilde{h}_{j}(\mathcal{Q})\}=E\{(\mathcal{Q}-m)\tilde{h}_{j}(\mathcal{Q})\}=0$ for all $j=1,\ldots,r_{\tilde{i}}$. Consequently, the tangent space $\mathcal{T}^{c}_{\bar{g}_{0}}$ of $\mathcal{P}^{c}_{\bar{g}}$ in (46) is given by:
    $\mathcal{T}^{c}_{\bar{g}_{0}}=\left\{l\in\mathcal{T}^{u}_{g_{0}}|E\{(\mathcal{Q}-m)l(\mathcal{Q})\}=0\right\}=\left\{l\in\mathcal{T}^{u}_{g_{0}}|E\{\mathcal{Q}l(\mathcal{Q})\}=0\right\}\subseteq\mathcal{H}.$ (191)
    We now need to derive the orthogonal projection operator onto $\mathcal{T}^{c}_{\bar{g}_{0}}$, i.e. $\Pi(\cdot|\mathcal{T}^{c}_{\bar{g}_{0}})$. This can be readily done by noticing that $\mathcal{T}^{u}_{g_{0}}$ can be decomposed as the following direct sum $\mathcal{T}^{u}_{g_{0}}=\mathcal{T}^{c}_{\bar{g}_{0}}\oplus\mathrm{Span}\{\mathcal{Q}-m\}$. This follows by the fact that: $E\{(\mathcal{Q}-m)l(\mathcal{Q})\}=0\Leftrightarrow(\mathcal{Q}-m)\perp l(\mathcal{Q})$ where we have used the canonical inner product of $\mathcal{H}$ defined in (2) as $\left\langle h_{1},h_{2}\right\rangle_{\mathcal{H}}\triangleq{E_{0}\{h_{1}h_{2}\}},\;\forall h_{1},h_{2}\in\mathcal{H}$. [^11] Note that we need to assume here the additional assumption $E\{\mathcal{Q}^{2}\}<\infty$. We note that this condition implies the assumption required in footnote 3. It imposes for example for the $t$-distribution that the degree of freedom satisfies $\nu>4$. Moreover, the previous decomposition of $\mathcal{T}^{u}_{g_{0}}$ implies that
    $\Pi(h|\mathcal{T}^{u}_{g_{0}})=\Pi(h|\mathcal{T}^{c}_{g_{0}})+\Pi(h|\mathrm{Span}\{(\mathcal{Q}-m)\}),\quad\forall h\in\mathcal{H},$ (192)
    where
[…]
    Finally, since $\Pi(h|\mathcal{T}^{u}_{g_{0}})=E\{h|\mathcal{Q}\}$, we have:
    $\Pi(h|\mathcal{T}^{c}_{\bar{g}_{0}})=E\{h|\mathcal{Q}\}-\frac{E\left\{\mathcal{Q}h\right\}}{E\left\{(\mathcal{Q}-m)^{2}\right\}}(\mathcal{Q}-m),\quad\forall h\in\mathcal{H},$ (194)
    where $E\{\mathcal{Q}h\}=E\{\mathcal{Q}E\{h|\mathcal{Q}\}\}=\left\langle\Pi(h|\mathcal{T}^{u}_{g_{0}}),(\mathcal{Q}-m)\right\rangle_{\mathcal{H}}$. To conclude, let us have a look at the nuisance tangent space of the parametric sub-model $\mathcal{P}^{c}_{\bm{\gamma},s,\bm{\zeta}}\subset\mathcal{P}_{\bm{\eta},\bar{g}}$ in (47). It is immediate to verify that its parametric sub-model $\mathcal{P}^{c}_{\bm{\gamma}_{0},s_{0},\bm{\zeta}}$ is exactly of the form given in (180). Consequently, its (finite-dimensional) tangent space $\mathcal{T}^{c}_{\bm{\zeta}_{0}}$ is included in $\mathcal{T}^{c}_{\bar{g}_{0}}$, i.e. $\mathcal{T}^{c}_{\bm{\zeta}_{0}}\subset\mathcal{T}^{c}_{\bar{g}_{0}}$.
  9. .9
    A property of the scale function $S(\bm{\Sigma}_{0})=|\bm{\Sigma}_{0}|^{1/m}$. The aim of this appendix is to recall the proof, already derived in [11], of the determinant-based scale. Specifically, as corollary of the Proposition IV.1, in [11], it has been show that, if the scale function $S_{d}(\bm{\Sigma}_{0})\triangleq|\bm{\Sigma}_{0}|^{1/m}$ is adopted, then the parametric FIM $\mathbf{I}_{\bm{\gamma}_{0}}$ in the parametric model $\mathcal{P}_{\bm{\gamma}}$ in (43) is equal to the efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})$ for $\mathcal{P}_{\bm{\eta},\bar{g}}$. To prove this, we need to show that, if $S(\bm{\Sigma}_{0})=|\bm{\Sigma}_{0}|^{1/m}$, then $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$. From (87), we have that:
    or equivalently $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$ iff $\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\in\text{Ker}\;\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})=\{a\cdot\mathrm{vec}(\mathbf{D}_{S}^{\mathbf{V}_{S,0}}),a\in\mathbb{R}\}$ where the matrix $\mathbf{D}_{S}^{\mathbf{V}_{S,0}}$ is formally defined in (175) of .7. As already noted in [11, Th. 3.1], this is a direct consequence of the Property P4 of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S,0}}$ (see appendix .7). In fact, for $S\equiv S_{d}$, we have that $\mathbf{D}_{S_{d}}^{\bm{\Sigma}_{0}}=m^{-1}|\bm{\Sigma}_{0}|^{1/m}\bm{\Sigma}_{0}^{-1}=m^{-1}\mathbf{V}_{S,0}^{-1}$. Consequently, we immediately have that $\mathrm{vec}(\mathbf{V}_{S_{d},0}^{-1})\in\text{Ker}\;\mathbf{M}_{S_{d}}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ and then $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$.
  10. .10
@@
    Explicit expressions of the CRB in the model $\mathcal{P}^{c}_{\bm{\gamma},s}$. Let us first evaluate the CRB on $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}\in\Omega$. To this end, we note that the two components of the related score vector $\mathbf{s}_{\bm{\nu}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T})^{T}$ have been already introduced in (56) and (57). Consequently, from standard calculation and by using the independence between $\mathcal{Q}$ and $\mathbf{u}$ (along with the properties of $\mathbf{u}$), the FIM for $\bm{\nu}_{0}\in\Omega$ is given by:
@@
    $\mathbf{I}_{\bm{\mu}_{0}}\triangleq E_{0}\left\{\mathbf{s}_{\bm{\mu}_{0}}\mathbf{s}_{\bm{\mu}_{0}}^{T}\right\}=\beta(\bar{g}_{0})\bm{\Sigma}_{0}^{-1},$ (66)
@@
    $\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\triangleq E_{0}\left\{\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T}\right\}=\mathbf{D}_{m}^{T}\left[\frac{1}{2}\alpha(\bar{g}_{0})(\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1})+\frac{1}{4}({\alpha(\bar{g}_{0})-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}\right]\mathbf{D}_{m}.$ (198)
    Then the CRB on $\bm{\nu}_{0}$ can be derived from the block diagonal structured FIM in (196) as:
@@
    $\mathrm{CRB}(\bm{\mu}_{0})\triangleq\mathbf{I}_{\bm{\mu}_{0}}^{-1}=\beta(\bar{g}_{0})^{-1}\bm{\Sigma}_{0}\ \ \mbox{and}\ \ \mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))\triangleq\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{-1}.$ (200)
@@
    $\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]^{-1}=2\alpha^{-1}(\bar{g}_{0})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{2\alpha^{-1}(\bar{g}_{0})(\alpha(\bar{g}_{0})-1)}{(m+2)\alpha(\bar{g}_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}.$ (201)
    Then using $\mathbf{D}_{m}\mathbf{D}_{m}^{\#}=\frac{1}{2}(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})$ and $\mathbf{K}_{m}(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})=(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})\mathbf{K}_{m}$ [25, Ch. 3], we straightforwardly get
    $\mathbf{D}_{m}\mathbf{D}_{m}^{\#}[\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]^{-1}\mathbf{D}_{m}^{\#T}\mathbf{D}^{T}_{m}=\alpha^{-1}(\bar{g}_{0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{2\alpha^{-1}(\bar{g}_{0})(\alpha(\bar{g}_{0})-1)}{(m+2)\alpha(\bar{g}_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}.$ (202)
    Finally, using $\mathbf{K}_{m}\mathbf{D}_{m}=\mathbf{D}_{m}$, $\mathbf{D}_{m}^{\#}\mathbf{D}_{m}=\mathbf{I}_{m(m+1)/2}$, [25, Ch. 3], it is easy to check that
@@
    and thus, since $\mathbf{D}_{m}^{\#}{\mathbf{K}}_{m}=\mathbf{D}_{m}^{\#}$ [25, Ch. 3]
    $\mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))=2\alpha^{-1}(\bar{g}_{0})\mathbf{D}_{m}^{\#}\left[(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{(\alpha(\bar{g}_{0})-1)}{(m+2)\alpha(\bar{g}_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}\right]\mathbf{D}_{m}^{\#T}.$ (204)
@@
    $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})=\alpha^{-1}(\bar{g}_{0})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},$ (213)